OpenStudy (anonymous):

Physics *Important question* One concave & one convex lens both having focal lengths 4cm are placed very closed to each other. Their combined focal length is: a) 2 cm b) 4cm c) 0 d) 1cm e) Infinity

OpenStudy (anonymous):

@ganeshie8 @mathslover @Mashy @secret66 @adrynicoleb @nincompoop @helpme1.2 @inkyvoyd @wolfe8 @Miracrown @ajeetkumarseo @Haseeb96

OpenStudy (anonymous):

@hartnn @iambatman

hartnn (hartnn):

if there is some distance 'd' between them, then we can use this formula Effective focal length \(EFD = \dfrac{f_1\times f_2}{f_1+f_2-d}\)

hartnn (hartnn):

in our case, f1 = 4, f2 = -4 d=0

hartnn (hartnn):

The first link gives you the derivation too :)

OpenStudy (anonymous):

let me solve it :)

OpenStudy (anonymous):

-16/0

hartnn (hartnn):

yeah, but i don't see \(\large -\infty \) as one of the choice.

OpenStudy (anonymous):

e is infinity

hartnn (hartnn):

but its definitely not a finite number, so i'd go with +infinity

OpenStudy (anonymous):

So E is answer .

OpenStudy (anonymous):

Right ? @hartnn

hartnn (hartnn):

yes, it is :)

OpenStudy (anonymous):

Thank you so much ^_^ @hartnn Chocolate cake for you :P

hartnn (hartnn):

yaay! nomnomnom :D welcome ^_^

OpenStudy (anonymous):

^_^